Please help me to solve this problem
Y=2sinx

Answer by Theo(13342)

you cannot solve this problem unless you know the value of x, or you know the value of y.

an example:

suppose x = 30 degrees.

the formula of y = 2 * sin(x) becomes y = 2 * sin(30).

you would use your calculator to find the value of sin(30), and you will get sin(30) = 1/2 = .5.

the formula then becomes y = 2 * .5 which will get you y = 1.

now, suppose you were given that y = 1.

the equation of y = 2sin(x) becomes 1 = 2sin(x).

you would divide both sides of the equaiton by 2 to get 1/2 = sin(x).

you would use the trigonometric rule that, if sin(x) = 1/2, then arcsin(1/2) = x.

you would use your calculator again to find that arcsin(1/2) = 30 degrees.

but, ........

unless you know the value of x, or the value of y, there's no way you can solve for the other.